# mycapequ[][List of equations]

### v1

@@ Abstract
Index Terms:
Hybrid Key Growing, Coherence Witness, Fock Space, Time Bin, Optical Fibers, Quantum Noise.
@@ I Introduction
Quantum Key Growing (QKG) [1, 2, 3, 4, 5, 6] is a cryptographic process that allows two parties—commonly referred to as Alice and Bob—to expand a small, initially shared secret key by exploiting the principles of quantum mechanics. More broadly, Quantum Key Distribution (QKD) refers to schemes that enable two parties to establish a shared secret key, even in the absence of an initial preshared secret. However, since QKD protocols often require a preshared key to ensure entity authentication, the term Quantum Key Growing more accurately captures the underlying process in most practical implementations. Current QKG protocols are subject to several stringent limitations, some of which may be intrinsic to the underlying physical systems [7, 8]. A prominent example is prepare-and-measure-based QKG [9, 10], which typically relies on quantum communication over optical fibers [11, 12]. These channels experience exponential signal attenuation with increasing transmission distance, thereby severely constraining the achievable range. To address this issue, satellite-based quantum links have been proposed as a scalable alternative [13, 14]. In contrast to fiber links, satellite channels exhibit signal loss that scales only quadratically with distance, making them more suitable for long-range quantum communication [15, 16]. However, their performance remains highly susceptible to atmospheric effects [17], including diurnal variations [18, 19], which in turn complicate reliable simulation and prediction [20]. Quantum relays offer a practical solution to the range limitation in QKG by forwarding signals through intermediate nodes [21, 22]. However, this approach requires that each relay node be fully trusted, thereby compromising the unconditional security often associated with QKG protocols. In contrast, quantum repeaters [23, 24, 25] enable long-distance, entanglement-based QKG [26, 27] without the need to trust intermediate nodes. Despite the fact that their practical deployment remains technologically challenging [28], the development of protocols designed for repeater-based architectures ensures a smoother transition to a future in which quantum and classical communication coexist [29, 30]. In parallel, prepare-and-measure QKG protocols—such as the one presented in this work—remain highly relevant due to their technological maturity and wide range of practical applications. These include secure short-distance communication (up to a few hundred kilometers) and scenarios in which the use of trusted quantum relays is considered acceptable.
@@
We introduce a novel method for Hybrid Key Growing (HKG) by combining quantum principles and one classical assumption on the physical layer—namely, that eavesdropping the quantum channel involved introduces a signal delay. Our protocol incorporates the main strength of one of the most practically robust variants of QKG to date: decoy-state QKG [10, 31, 32]. Like the decoy-state approach, our protocol does not rely on ideal single-photon sources or detectors, thereby improving resilience to moderate photon loss and enhancing practical security. A central advantage of our method is that it eliminates the basis reconciliation (key-sifting) step found in all conventional QKG protocols, effectively doubling the bit-per-pulse key generation rate. Existing schemes that also utilize multiple photonic Degrees of Freedom (DoFs) [33] can achieve similar benefits, but typically incur a higher classical communication overhead. Our protocol, by contrast, adds complexity primarily at the transmitter and receiver hardware level, leaving its impact on the overall bit-per-second rate uncertain. Nevertheless, if this complexity does not significantly hinder throughput, the resulting gains could be highly valuable in latency-sensitive applications such as Digital Twin (DT) systems [34, 35], where the rapid establishment and renewal of shared secret keys is essential. Our HKG protocol introduces a novel approach to handling quantum noise by leveraging full knowledge of the underlying two-qubit quantum channel—a reasonable assumption in many practical scenarios [36, 37]. We optimize $L$, a defining parameter of our protocol, to maximize performance under given noise conditions. Our simulations reveal that, under low-noise conditions, our encoding strategy decreases the Quantum Bit Error Rate (QBER) without compromising security.
@@ II-A Optical fibers: loss, dephasing and dark counts
where the underlying Hilbert space is referred to the so-called Fock basis and $p_{\eta}(k,M)\equiv\binom{M}{k}(1-\eta)^{k}\eta^{M-k}$, with $\eta=10^{-\frac{d\alpha}{10}}$ and $\alpha\approx 2\cdot 10^{-1}\hskip 2.84544pt\frac{\mathrm{dB}}{\mathrm{km}}$ [28]. On the other hand, the dephasing channel $\Lambda_{\gamma}$ can be described as [39]
@@
where $\rho$ is a general state, $f_{\gamma}(x)\equiv e^{-\frac{\gamma}{2}x^{2}}$, and the dephasing parameter $\gamma$ scales linearly with time [40] and, hence, with $d$.
@@
where $p_{\lambda}(k)\equiv\frac{\lambda^{k}e^{-\lambda}}{k!}$, with $\lambda\approx 10^{-3}$ [41] being the average number of dark counts within the selected time window.
@@ II-B Bosonic encoding: a way out to moderate photon loss
leading to an enhanced readout distinguishability by assigning a ”0” to all photon counts $m$ fulfilling
@@ II-C Coherence witness
Given the dephasing operation $\Delta$[43] and any Hermitian operator $\sigma$, we define a coherence witness as
$W_{\sigma}\equiv\Delta(\sigma)-\sigma.$ (10)
As shown in [43], if the expectation of $W_{\sigma}$ over a certain state is negative, then that state must exhibit quantum coherence, or as often also termed: quantum superposition.
@@
If $\sigma$ is a density matrix, the expectation of $W_{\sigma}$ over an ensemble of quantum states can be computed using the swap test [44, 45].
@@ III A new HKG protocol
In the first phase, Alice, the sender, samples a bitstring $\mathbf{K}$ of length $\beta=\nu\omega$, with $\omega,$ $\nu\in\mathbb{N}$ and with $\nu>\omega$. Subsequently, she encodes each bit of $\mathbf{K}$ as indicated by another bitstring $\mathbf{F}$, of the same length and preshared with Bob:
If a certain bit of $\mathbf{F}$ is a ”0” Alice encodes the corresponding bit $b_{0}$ of $\mathbf{K}$ as
@@
where kets with the subscript time-bin indicate qubit states belonging to the Hilbert space spanned by two distinct and orthogonal pulse time bins, and $\xi$ is a uniformly random bit Alice needs to sample at each preparation. Instead, if a certain bit of $\mathbf{F}$ is a ”11”, Alice encodes the corresponding bit $b_{1}$ of $\mathbf{K}$ as
@@
Bob receives the sent pulses, and performs different operations on them, depending on their type. He measures those prepared as in Eq. (11) in the Fock basis to extract, for each of them, a secret bit $\hat{b}_{0}$. Given low quantum noise, the bit $\hat{b}_{0}$ likely coincides with the bit $b_{0}$ encoded by Alice. Moreover, the time-bin Degree of Freedom (DoF) of such incoming pulses is coupled with another reference system prepared as
@@
$|W_{\mathrm{time\text{-}bin}}\equiv W_{\ket{\phi}\bra{\phi}}$ (14)
over certain input states. On the other hand, Bob measures the time-bin of the incoming pulses prepared as in Eq. (12) to obtain, for each pulse, an estimated bit $\hat{b}_{1}$. Being $L$ a natural parameter chosen after channel calibration (see Section V-A), Bob interacts with the photon-number property of each incoming pulse by coupling it with another reference system
@@
Alice and Bob publicly compare a small random fraction $\kappa$, discarded for later use, of the obtained raw key. A Quantum Bit Error Rate (QBER) smaller than a threshold $\mu<\frac{1}{4}$, defined during channel calibration, leads Alice to reckon the receiver as legitimate, i.e., as Bob. Otherwise, the protocol is aborted.
@@
$\langle W_{\mathrm{time\text{-}bin}}\rangle_{\mathrm{selection_{1}}}\hskip 5.69046pt\mathrm{and}\hskip 5.69046pt\langle W_{\mathrm{Fock}}(L)\rangle_{\mathrm{selection_{2}}},$ (17)
and computes the average $\hat{W}(L)$ between them. An obtained value below a threshold $\frac{1}{2}<\tau<0$, defined during channel calibration, ensures both no eavesdropping and sender’s legitimacy. Otherwise, the protocol is aborted.
@@
Our protocol does not require the commonly used bases-reconciliation step, thus reducing the classical communication overhead and approximately doubling, due to no bases mismatching, the bit-per-pulse key rate with respect to conventional QKG schemes.
@@
$\mathbf{F}$ is divided into $\omega$ segments, each of $\nu$ bits by construction. Alice and Bob select, and discard for later use, $\omega$ of the secret shared bits generated. For bits of value ”11”, their corresponding segments in $\mathbf{F}$ are flipped. Otherwise, i.e., for bits of value ”0”, the corresponding bit segments remain unchanged. Furthermore, in any $k$-th round, with $k>1$, Bob’s authentication (see 3rd phase) takes place by him and Alice comparing $\omega$ sets of bits, each set belonging to one of the defined segments, and each set being of size $\kappa\nu$. Bob is accepted if the QBER is smaller than the threshold $\mu$ in all $\omega$ instances and, otherwise, the protocol is aborted. Under no protocol abortion, Bob confirms Alice’s identity by computing $\hat{W}$ for each corresponding set. Alice is accepted if $\hat{W}<\tau$ in all $\omega$ instances and, otherwise, the protocol is aborted.
@@ IV-A Authentication completeness
With regard to completeness, Bob is trivially accepted given no malicious parties and no quantum noise. In such case, and setting $L=N$, Alice is also accepted, because Bob performs the swap test against a state that flags each coherence witness in all instances. Namely, idealistically [50, 51],
$\langle W_{\mathrm{time\text{-}bin}}\rangle_{\mathrm{selection_{1}}}=-\frac{1}{2}\hskip 5.69046pt\mathrm{and}$ (18)
$\small\langle W_{\mathrm{Fock}}(N)\rangle_{\mathrm{selection_{2}}}=-\frac{1}{2}.$ (19)
Moreover, given the continuity of both the expected QBER and $\hat{W}$ with respect to state fidelity, an adequate choice of $\mu$ and $\tau$ ensures completeness in setups with low-noise conditions.
@@ IV-B1 Resistance against Bob-impersonation attacks
Firstly, let us consider an adversary not knowing $\mathbf{F}$, and performing a straightforward man-in-the-middle attack [52]. For each incoming pulse, the probability of measuring the right DoF is $\frac{1}{2}$, thus leading to an expected QBER lower bounded by $\frac{1}{4}>\mu$. Secondly, notice that only eavesdropping the used quantum channel provides no information about $\mathbf{F}$. However, this combined with the disclosure of $\xi$-choices and the information reconciliation scheme–both of which are public by design–may compromise the secrecy of $\mathbf{F}$. Desirably, the 5th phase of our scheme ensures that, even with full knowledge of $\mathbf{F}$, the probability of correctly forging Bob is roughly lower than $\frac{1}{2^{\omega}}$. That is, an attacker knowing $\mathbf{F}$, prior to its update, and aiming to impersonate Bob, measures with a high probability $p>1-\frac{1}{2^{\omega}}$ the wrong DoF for, at least, one complete subset of $\nu$ pulses. Hence, the expected QBER for the corresponding segment is $\frac{1}{2}>\mu$.
@@ IV-B2 Resistance against Alice-impersonation attacks
Two conditions are needed to impersonate Alice: after sampling $\beta$ bits, encoding each of them in a light pulse and sending them to Bob, both the QBER and $\hat{W}$ obtained must be low enough. This renders the attacker’s encoding choices nontrivial. If only one condition was to be fulfilled, an adversary could double-encode a bitstring (to avoid a high QBER) or double-encode superposition states (to avoid a high value of $\hat{W}$). We again first consider an adversary with no knowledge of $\mathbf{F}$. Each of the adversary’s encoding guesses is successful with a probability $\frac{1}{2}$. Therefore, on average, half of the times in which Bob would perform the swap test against a state flagging a coherence witness, he does so with non-superposition states, thus substantially increasing $\hat{W}$. Moreover, such attack would induce an expected QBER of $\frac{1}{4}$. Finally, an attacker knowing $\mathbf{F}$, and making a guess on the key update, encodes with high probability $p>1-\frac{1}{2^{\omega}}$ all $\nu$ bits of at least one of the $\omega$ underlying subsets in the wrong DoF. In such case, the expected QBER for at least one of the bit-set compared rises to $\frac{1}{2}$. Moreover, the value of $\hat{W}$ would vanish in those sets of bits corresponding to a wrong $F$-segment update.
@@ IV-C Key secrecy
As stated above, $\mathbf{F}$ may be leaked to an eavesdropper harnessing public discussions and measurements on the sent stream of light pulses. This can compromise key secrecy against the well-known intercept-resend [53] attack. However, we make an assumption on the physical layer: an intercept-resend attack would introduce substantial time delays on the light pulses, thereby perturbing their time-bin DoF and thus increasing both the QBER and the value of $\hat{W}$. This assumption, which does not rely on the laws of quantum mechanics, but rather on the specifics of the channel engineering, is what renders our protocol hybrid.

### v2

@@ I Introduction
Quantum Key Growing (QKG) [1, 2, 3] is a cryptographic process that allows two parties—commonly referred to as Alice and Bob—to expand a small, initially shared secret key by exploiting the principles of quantum mechanics. More broadly, Quantum Key Distribution (QKD) refers to schemes that enable two parties to establish a shared secret key, even in the absence of an initial preshared secret. However, since QKD protocols often require a preshared key to ensure entity authentication, the term Quantum Key Growing more accurately captures the underlying process in most practical implementations. Current QKG protocols face several stringent limitations, some of which may be inherent to the underlying physical systems [4, 5]. A notable example arises in prepare-and-measure-based QKG [6, 7], which typically relies on quantum communication through optical fibers [8, 9]. These channels suffer from exponential signal attenuation with increasing distance, significantly limiting the achievable transmission range. To address this issue, satellite-based quantum links have been proposed as a scalable alternative [10, 11]. In contrast to fiber links, satellite channels exhibit signal loss that scales only quadratically with distance, making them more suitable for long-range quantum communication [12, 13]. However, their performance remains highly susceptible to atmospheric effects [14], including diurnal variations [15, 16], which in turn complicate reliable simulation and prediction [17]. Quantum relays offer a practical solution to the range limitation in QKG by forwarding signals through intermediate nodes [18, 19]. However, this approach requires complete trust in each relay node, which is generally undesirable. In contrast, quantum repeaters [20, 21, 22] enable long-distance, entanglement-based QKG [23, 24] without the need to trust intermediate nodes. Despite the fact that their practical deployment remains technologically challenging [25], the development of protocols designed for repeater-based architectures ensures a smoother transition to a future in which quantum and classical communication coexist [26, 27]. In parallel, prepare-and-measure QKG protocols remain highly relevant due to their technological maturity and wide range of practical applications. These include secure short-distance communication (up to a few hundred kilometers) and scenarios in which the use of trusted quantum relays is considered acceptable.
@@
We propose a novel Hybrid Key Growing (HKG) method that combines quantum principles—drawing from prepare-and-measure schemes—with a single classical assumption: that eavesdropping on the quantum channel causes a detectable delay. Our protocol leverages the main strength of one of the most practical prepare-and-measure QKG variants: the decoy-state technique [7, 28, 29]. Like the decoy-state approach, our protocol does not rely on ideal single-photon sources or detectors, thereby improving resilience to moderate photon loss and enhancing practical security. A central advantage of our method is that it eliminates the basis reconciliation (key-sifting) step found in all conventional QKG protocols, effectively doubling the bit-per-pulse key generation rate. Existing schemes that also utilize multiple photonic Degrees of Freedom (DoFs) [30] can achieve similar benefits, but typically incur a higher classical communication overhead. Our protocol, by contrast, adds complexity primarily at the transmitter and receiver hardware level, leaving its impact on the overall bit-per-second rate uncertain. Nevertheless, if this complexity does not significantly hinder throughput, the resulting gains could be highly valuable in latency-sensitive applications such as Digital Twin (DT) systems [31, 32], where the rapid establishment and renewal of shared secret keys is essential in security-sensitive applications. Our HKG protocol introduces a novel approach to handling quantum noise by leveraging full knowledge of the underlying two-qubit quantum channel—a reasonable assumption in many practical scenarios [33, 34]. We optimize $L$, a defining parameter of our protocol, to maximize performance under given noise conditions. Our simulations reveal that, under low-noise conditions, our encoding strategy decreases the Quantum Bit Error Rate (QBER) without compromising security.
@@ II-A Optical fibers: loss, dephasing and dark counts
where the underlying Hilbert space is referred to the so-called Fock basis and $p_{\eta}(k,M):=\binom{M}{k}(1-\eta)^{k}\eta^{M-k}$, with $\eta=10^{-\frac{d\alpha}{10}}$ and $\alpha\approx 2\cdot 10^{-1}\hskip 2.84544pt\frac{\mathrm{dB}}{\mathrm{km}}$ [25]. On the other hand, the dephasing channel $\Lambda_{\gamma}$ can be described as [36]
@@
where $\rho$ is a general state, $f_{\gamma}(x):=e^{-\frac{\gamma}{2}x^{2}}$, and the dephasing parameter $\gamma$ scales linearly with time [37] and, hence, with $d$.
@@
where $p_{\lambda}(k):=\frac{\lambda^{k}e^{-\lambda}}{k!}$, with $\lambda\approx 10^{-3}$ [38] being the average number of dark counts within the selected time window.
@@ II-B Bosonic encoding: a way out to moderate photon loss
leading to an enhanced readout distinguishability by assigning a “0” to all photon counts $m$ fulfilling
@@ II-C Coherence witness
Given the dephasing operation $\Delta$[40] and any Hermitian operator $\sigma$, we define a coherence witness as the observable
$\hat{W}_{\sigma}:=\Delta(\sigma)-\sigma.$ (10)
As shown in [40], if the expectation of $\hat{W}_{\sigma}$ over a certain state is negative, then that state must exhibit quantum coherence, or as often also termed: quantum superposition.
@@
If $\sigma$ is a density matrix, the expectation of $\hat{W}_{\sigma}$ over an ensemble of quantum states can be computed using the swap test [41, 42, 43].
@@ III A new HKG protocol
In the first phase, Alice, the sender, generates a bitstring $\mathbf{K}$ uniformly at random of length $\beta=\nu\omega$, with $\omega,$ $\nu\in\mathbb{N}$ and with $\nu>\omega$. Subsequently, she encodes each bit of $\mathbf{K}$ as indicated by another bitstring $\mathbf{F}$, also generated uniformly at random, of the same length as $\mathbf{K}$, and preshared with Bob: If a certain bit of $\mathbf{F}$ is a “0” Alice encodes the corresponding bit $b_{0}$ of $\mathbf{K}$ as
@@
where kets with the subscript time-bin indicate qubit states belonging to the Hilbert space spanned by two distinct and orthogonal pulse time bins, and $\xi$ is a uniformly random bit Alice needs to generate uniformly at random at each preparation. Instead, if a certain bit of $\mathbf{F}$ is a “11”, Alice encodes the corresponding bit $b_{1}$ of $\mathbf{K}$ as
@@
Bob receives the sent pulses, and performs different operations on them, depending on their type. He measures those prepared as in Eq. (11) in the Fock basis [44] to extract, for each of them, a secret bit $\tilde{b}_{0}$. Given low quantum noise, the bit $\tilde{b}_{0}$ likely coincides with the bit $b_{0}$ encoded by Alice. Moreover, the time-bin Degree of Freedom (DoF) of such incoming pulses is coupled with another reference system prepared as
@@
$|\hat{W}_{\mathrm{time\text{-}bin}}:=\hat{W}_{\ket{\phi}\bra{\phi}}$ (14)
over certain input states. On the other hand, Bob measures the time-bin of the incoming pulses prepared as in Eq. (12) to obtain, for each pulse, an estimated bit $\tilde{b}_{1}$. Being $L$ a natural parameter chosen after channel characterization (see Section V-A), Bob interacts with the photon-number property of each incoming pulse by coupling it with another reference system
@@
Alice and Bob publicly compare a small fraction $\kappa$, discarded for later use, of the obtained raw key. A Quantum Bit Error Rate (QBER) smaller than a threshold $\mu<\frac{1}{4}$, defined during channel characterization, leads Alice to regard the receiver as legitimate. Otherwise, the protocol is aborted.
@@
$\langle\hat{W}_{\mathrm{time\text{-}bin}}\rangle_{\mathrm{selection_{1}}}\hskip 5.69046pt\mathrm{and}\hskip 5.69046pt\langle\hat{W}_{\mathrm{Fock}}(L)\rangle_{\mathrm{selection_{2}}},$ (17)
and computes the average $W(L)$ between them. An obtained value below a threshold $\frac{1}{2}<\tau<0$, defined during channel characterization, ensures both no eavesdropping and sender’s legitimacy. Otherwise, the protocol is aborted.
@@
Our HKG protocol does not require the commonly used bases-reconciliation step, thus reducing the classical communication overhead and approximately doubling, due to no bases mismatching, the bit-per-pulse key rate with respect to conventional QKG schemes.
@@
$\mathbf{F}$ is divided into $\omega$ segments, each of $\nu$ bits by construction. Alice and Bob select, and discard for later use, $\omega$ of the secret shared bits generated. For bits of value “11”, their corresponding segments in $\mathbf{F}$ are flipped. Otherwise, i.e., for bits of value “0”, the corresponding bit segments remain unchanged. Furthermore, in any $k$-th round, with $k>1$, Bob’s authentication (see 3rd phase) takes place by him and Alice comparing $\omega$ sets of bits, each set belonging to one of the defined segments, and each set being of size $\kappa\nu$. Bob is accepted if the QBER is smaller than the threshold $\mu$ in all $\omega$ instances and, otherwise, the protocol is aborted. Under no protocol abortion, Bob confirms Alice’s identity by computing $W$ for each corresponding set. Alice is accepted if $W<\tau$ in all $\omega$ instances and, otherwise, the protocol is aborted.
@@ IV-A Authentication completeness
With regard to completeness, Bob is trivially accepted given no malicious parties and no quantum noise. In such a case, and setting $L=N$, Alice is also accepted, because Bob performs the swap test against a state that flags each coherence witness in all instances. Namely, ideally [49, 50],
$\langle\hat{W}_{\mathrm{time\text{-}bin}}\rangle_{\mathrm{selection_{1}}}=-\frac{1}{2}\hskip 5.69046pt\mathrm{and}$ (18)
$\small\langle\hat{W}_{\mathrm{Fock}}(N)\rangle_{\mathrm{selection_{2}}}=-\frac{1}{2}.$ (19)
Moreover, given the continuity of both the expected QBER and $W$ with respect to quantum-state fidelity, an adequate choice of $\mu$ and $\tau$ ensures completeness in setups with low-noise conditions.
@@ IV-B1 Resistance against Bob-impersonation attacks
Firstly, let us consider an adversary not knowing $\mathbf{F}$, and performing a straightforward man-in-the-middle attack [51]. For each incoming pulse, the probability of measuring the right DoF is $\frac{1}{2}$, thus leading to an expected QBER lower bounded by $\frac{1}{4}>\mu$. Secondly, notice that only eavesdropping the used quantum channel provides no information about $\mathbf{F}$. However, such an eavesdropping combined with the disclosure of the choices of each $\xi$ and the information reconciliation scheme—both of which are public by design—may compromise the secrecy of $\mathbf{F}$. Desirably, the 5th phase of our scheme ensures that, even with full knowledge of $\mathbf{F}$, the probability of correctly forging Bob is roughly equal to $\frac{1}{2^{\omega}}$. That is, an attacker knowing $\mathbf{F}$, prior to its update, and aiming to impersonate Bob, measures with a high probability $p=1-\frac{1}{2^{\omega}}$ the wrong DoF for, at least, one complete subset of $\nu$ pulses. Hence, the expected QBER for such a subset is $\frac{1}{2}>\mu$.
@@ IV-B2 Resistance against Alice-impersonation attacks
Two conditions are needed to impersonate Alice: after generating $\beta$ bits, encoding each of them in a light pulse and sending them to Bob, both the QBER and $W$ obtained must be low enough. This renders the attacker’s encoding choices nontrivial. If only one condition was to be fulfilled, an adversary could double-encode a bitstring (to avoid a high QBER) or double-encode superposition states (to avoid a high value of $W$). We again first consider an adversary with no knowledge of $\mathbf{F}$. Each of the adversary’s encoding guesses is successful with a probability $\frac{1}{2}$. Therefore, on average, half of the times in which Bob, under no forgery attempt, would perform the swap test against a state flagging a coherence witness, he does so with non-superposition states, thus substantially increasing $W$. Moreover, such an attack would induce an expected QBER of $\frac{1}{4}$. Finally, an attacker who knows $\mathbf{F}$, and makes a guess about the key update, encodes with high probability $p=1-\frac{1}{2^{\omega}}$ all $\nu$ bits of at least one of the $\omega$ underlying subsets in the wrong DoF. In this case, the QBER in at least one compared bit set would be $\frac{1}{2}$, and the value of $W$ in this set would drop to zero.
@@ IV-C Key secrecy
As stated above, $\mathbf{F}$ may be leaked to an eavesdropper harnessing public discussions and measurements on the sent stream of light pulses. This can compromise key secrecy against the well-known intercept-resend [52] attack. However, we make an assumption about the physical layer: an intercept-resend attack would introduce substantial time delays on the light pulses, thus perturbing their time-bin DoF and thus increasing both the QBER and the value of $W$. This assumption, which relies on the specifics of the channel engineering and the adversary’s classical capabilities, is what renders our protocol hybrid.